OpenStudy (anonymous):

Anyone know the easiest way to factor polynomials e.g. (x^2 + x^3)^3?

OpenStudy (anonymous):

know the rules maam

hero (hero):

I'll take a stab at it

hero (hero):

\((x^2 + x^3)^3 = (x^2 + x^3)(x^2 + x^3)(x^2 + x^3)\) \(=x^2(1 + x)x^2(1 + x)x^2(1+x)\) \(=(x^2)^3(1+x)^3\) \(=x^6(1+x)^3\) That last expression represents the factorization as a product of two factors.

OpenStudy (anonymous):

its true you could just simply expand which is more accurate from where you have x^6+3^7+3x^8+x^9 were you can easily see that x^6 is a factor already and so you could divide the polynomial by x^6 to get other factors.

OpenStudy (anonymous):

this method is just the use of simple expansion but hero's method is direct factorisation which is easier and faster

OpenStudy (anonymous):

Thank you Hero & nelotsak for your help!

hero (hero):

You're welcome

OpenStudy (anonymous):

I'm really struggling with Calculus right now which I last took in 1998 and have only used a very little bit since then. I am struggling with remembering/understanding the use of derivatives in Calculus which was part of that question I asked. The whole question is actually: Calculate y' if \[y = (x^2 + x^3 )^4\]

hero (hero):

Yes, it would be easier to find the derivative of \(y = (x^2 + x^3)^4\)

OpenStudy (anonymous):

So I'm gonna try to follow your method you showed me to get the final answer of: \[y \prime = 4x ^{7}(x+1)^{3}(3x+2)\]

hero (hero):

Good luck with it. You already have the answer so...

OpenStudy (anonymous):

I understand that the derivative of \[x ^{2}+x ^{3} = 3x ^{2}+2x\] but am confused to as to whether or not to factor the 3x & 2x out first or to the forth power. Thank you, again!

hero (hero):

Well if you factored out the x you would have \(x(3x + 2)\) which gets you closer to what you want.

OpenStudy (anonymous):

Cool. Thank you!

hero (hero):

So do you see how to get to the derivative now?

OpenStudy (anonymous):

Well when I first started to answer it I get \[4(3x^2 + 2x)^3\] but what I'm wondering is if I'm missing \[(x^2 + x^3)^3 and then finish factoring and solving \it as you showed me earlier\]

OpenStudy (anonymous):

That was supposed to say: and then finish factoring and solving as you showed me earlier.

hero (hero):

I messed up on one of the steps. I put an equal sign instead of a plus.

hero (hero):

I hate that because it is not easy to re-write

OpenStudy (anonymous):

Okay, I see it. I was confused how they were set equal to each other

hero (hero):

Okay, good. Long as you understand it.

OpenStudy (anonymous):

I do. I broke it down a little different and I'm not sure if it is correct: \[= 4 ( 3x^2 + 2x) (x^2 + x^3) ^3\] \[= 4x (3x + 2) (x^2)^3 (x +1)^3\] \[= 4x(x^6)(3x+2)(x+1)^3 = 4x^7(3x+2)(x+1)^3\]

hero (hero):

The difference is I started from the very beginning.

hero (hero):

I messed up twice. I should never have had a plus symbol in certain spots.

hero (hero):

I'm just going to delete what I wrote.

OpenStudy (anonymous):

No need. I understood what you you meant.

OpenStudy (anonymous):

So how would you do it differently if you need to solve for y' but are unable to factor y completely out of the equation e.g. \[xy^4 + x^2y = x + 3y\]

hero (hero):

You have to do implicit differentiation.

OpenStudy (anonymous):

Also, Hero, do you know any general rules of thumb of when to use the sum, scalar multiple, product, quotient, chain, inverse or implicit rule?

hero (hero):

Where is your textbook?

hero (hero):

All those rules should be in it.

OpenStudy (anonymous):

Right in front of me confusing the heck out of me. According to the math placement exam I was ready for Calculus 1 and understood it for a second but now I am LOST!

hero (hero):

Which textbook do you have?

OpenStudy (anonymous):

I'm looking at them right now. I'm using Calculus 7th ed. by James Stewart, McMaster University & University of Toronto

hero (hero):

I see. Trust me, that's an easy book compared to the book I had to use when I was taking Calculus.

OpenStudy (anonymous):

Good to hear...I think ;-)

hero (hero):

If you want help with calculus, you should try asking some of the other 99 users like wio, king george, satellite73. Ask wio for help.

OpenStudy (anonymous):

Okay, thank you. I have been meaning to look into getting a tutor to try to dumb it down for me so that I can start to understand again. I will take your advice. Thank you you again for your help and your time, Hero!

hero (hero):

You're welcome.

OpenStudy (anonymous):

Btw, what are 99 users in comparison to me, a "15 user"?

hero (hero):

A 99 user has quite a bit of experience on OS. Has helped a lot of users, answered (or asked) many questions and has a long tenure on OS. Most 99+ users have been here for at least a year. A few have been here for just six months. I came here before they started doing smart score, so when they converted to it, I was already 99 user.

OpenStudy (anonymous):

Cool, so I just need to lookup Wio and message him/her to see if they might possibly have a bit of time to help teach me something? :-D

hero (hero):

Maybe. You have to catch the person at a good time.

hero (hero):

If you show that you are willing to participate in the process and that you have a clue what's going on, then you're more likely to get helped.

OpenStudy (anonymous):

Thank you. I'll try not to appear as lost but unfortunately that is sort of where I am.

OpenStudy (anonymous):

USe the binomial theorem